OpenStudy (anonymous):

stepwise solution lim(--> infinity) (2x^2-1)^1/2-(2x^2-3x)^1/2

OpenStudy (turingtest):

whenever you wee something that is the difference of two square roots, you usually have to multiply by its conjugate

OpenStudy (turingtest):

see*

OpenStudy (turingtest):

\[\sqrt{2x^2-1}-\sqrt{2x^2-3x}\cdot{\sqrt{2x^2-1}+\sqrt{2x^2-3x}\over\sqrt{2x^2-1}+\sqrt{2x^2-3x}}\]

OpenStudy (turingtest):

remember that\[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

yeah. you're really good with math

OpenStudy (turingtest):

thanks, practice is all...

OpenStudy (anonymous):

ok, let me work this problem out and see if i get it

OpenStudy (turingtest):

I haven't tried it either, I just know how these usually go

OpenStudy (turingtest):

but smart people have given me medals, so I bet it's gonna work :)

OpenStudy (anonymous):

lol. you're smart too

OpenStudy (turingtest):

lol, only when I'm not busy being stupid 'gotta go, good luck!

OpenStudy (anonymous):

thank u! bye and take care

OpenStudy (anonymous):

actually this still stays \[\LARGE \infty\] I was trying to solve it, I guess I couldn't! just got a circle back to the beginning... L'Hospitals doesn't work either ... wolframa proves it ! http://www.wolframalpha.com/input/?i=lim_%7Bx%5Cto+%5Cinfty%7D+%282x%5E2-1%29%5E%281%2F2%29-%282x%5E2-3x%29%5E%281%2F2%29

OpenStudy (anonymous):

i couldn't solve it either. i thought it DNE

OpenStudy (anonymous):

DNE= ? :O

OpenStudy (anonymous):

does not exist

OpenStudy (anonymous):

ahh .. hahaha lol :P These shortcuts you use, I don't get them at all lol. :)

OpenStudy (anonymous):

lol. i used that wolfram website too and the answer was definite http://www.wolframalpha.com/input/?i=limits&a=*C.limits-_*Calculator.dflt-&f2= √%282x%5E2-1%29-√%282x%5E2-3x%29&x=4&y=8&f=Limit.limitfunction_√%282x%5E2-1%29-√%282x%5E2-3x%29&f3=∞&f=Limit.limit_∞&a=*FVarOpt.1-_**-.***Limit.limitvariable--.**Limit.direction--.**Limit.limitvariable2-.*Limit.limit2-.*Limit.direction2---.*--

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \left( \sqrt{2x^{2} -1}-\sqrt{2x^2 - 3x} \right)\]multiply by its conjugate\[\lim_{x \rightarrow \infty} \left( \sqrt{2x^{2} -1}-\sqrt{2x^2 - 3x} \right)\left( \frac{\sqrt{2x^2 - 1}+\sqrt{2x^2 - 3x}}{\sqrt{2x^2 - 1}+\sqrt{2x^2 - 3x}} \right)\]\[\lim_{x \rightarrow \infty}\left( \frac{3x - 1}{\sqrt{2x^2 - 1}+\sqrt{2x^2 - 3x}} \right)\left( \frac{\frac{1}{x}}{\frac{1}{x}} \right)\]\[\lim_{x \rightarrow \infty}\left( \frac{3-\frac{1}{x}}{\sqrt{2-\frac{1}{x^2}}+\sqrt{2-\frac{3}{x}}} \right)\]\[\frac{3-0}{\sqrt{2-0}+\sqrt{2-0}}\]\[\frac{3}{2\sqrt{2}}\]

OpenStudy (anonymous):

where did the 1/x come from?

OpenStudy (anonymous):

I multiply the numerator and denominator by 1/x

OpenStudy (anonymous):

oh ok. why?

OpenStudy (anonymous):

i think i know why. to cancel out the x from 3x?

OpenStudy (anonymous):

i don't understand why 1/x is 0 as x approaches \[\infty\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\frac{1}{x}=0\]the limit as x approaches infinity of (1/x) is 0, to understand this, try for a really large value for x, you will get results that are nearly 0

OpenStudy (anonymous):

oh ok! thank you. Also, when you multiplied by 1/x how did you derive the denominator to be \[\sqrt{2-1/x^2}\]

OpenStudy (anonymous):

\[\frac{\sqrt{2x^2 - 1}}{x}=\sqrt{\frac{2x^2 - 1}{x^2}}=\sqrt{2 - \frac{1}{x^2}}\]

OpenStudy (anonymous):

speachless... I feel soo stupid! :$ I should've realized the part when we have to divide by x but I only thought about substituting. I guess I got blind totally. Great job @exraven